Frank is 6 years older than Mark. three years ago Frank was 4 times as old as Mark, How old is Mark?
Answer by stanbon(75887)

Frank is 6 years older than Mark. three years ago Frank was 4 times as old as Mark, How old is Mark?
-----
Equations:
f = m + 6
(f-3) = 4(m-3)
---
Substitute for "f" and solve for "m":
(m+6-3) = 4m-12
m+3 = 4m-12
3m = 15
mark = 5 yrs
====================
cheers,
Stan H.
================